# The $u$-invariant of function fields in one variable

### v1

@@ 3. $(c)$
$(b)$ If $\varphi$ is isotropic, then so is $\overline{\varphi}^{v}$, by 3.8. If $\overline{\varphi}$ is isotropic, then $\overline{\varphi}^{\prime}=\langle\overline{a}_{2},\dots,\overline{a}_{n}\rangle_{Kv}$ represents $-\overline{a}_{1}$, and it follows by $(a)$ that $\varphi^{\prime}$ represents $-a_{1}$, whereby $\varphi$ is isotropic.
$(c)$ We prove the equivalence by induction on $n$. If $n=0$, then $\varphi$ and $\overline{\varphi}^{v}$ are trivially hyperbolic. Assume that $n>0$. If $\varphi$ is anisotropic, then so is $\overline{\varphi}^{\prime}$, by $(b)$, and as $n>0$, it follows that neither $\varphi$ nor $\overline{\varphi}$ is hyperbolic. Assume that $\varphi$ is isotropic. By 3.5 $(b)$, there exist $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\varphi\simeq\mathbb{H}_{K}\perp\langle b_{3},\dots,b_{n}\rangle_{K}$ and $\overline{\varphi}\simeq\mathbb{H}_{Kv}\perp\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$. Now, by Witt’s Cancellation Theorem [17, Chap. I, Theorem 4.2], $\varphi$ is hyperbolic if and only if $\rho=\langle b_{3},\dots,b_{n}\rangle_{K}$ is hyperbolic, and $\overline{\varphi}^{v}$ is hyperbolic if and only if $\overline{\rho}^{v}=\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$ is hyperbolic. By the induction hypothesis, $\rho$ is hyperbolic if and only if $\overline{\rho}^{v}$ is hyperbolic. From this together we conclude that $\varphi$ is hyperbolic if and only if $\overline{\varphi}^{v}$ is hyperbolic. ∎
@@
Part of the following statement recovers [9, Theorem 1] and its corollary.
@@
[…]
@@
When restricting attention to complete discrete valuation rings and their fraction fields, most (if not all) of the results in this section are well-known. To the best of our knowledge these first appeared explicitly in Durfee’s doctoral research [9] for fields with a complete rank $1$ valuation. Inspection reveals that Durfee’s proofs (and other versions found later) mostly work without changes for general valued fields $(K,v)$ with $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ and $2\in\mathcal{O}_{v}^{\times}$, thus in particular for non-dyadic henselian valued fields, which would be sufficient for the purposes in the rest of this article. For our own and the reader’s convenience, we have included here an elementary and self-contained exposition which postpones adding the assumption $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ until the very end.
@@ 4. The generalised $u$-invariant
Let $K$ be a field with $\mathsf{char}(K)\neq 2$. Recall that we defined a quadratic form $\varphi$ over the field $K$ to be a torsion from if $m\times\varphi=\varphi\perp\ldots\perp\varphi$ is hyperbolic for some $m\in\mathbb{N}^{+}$. The general $u$-invariant of $K$ as defined by Elman and Lam in [11] is given by
@@
Note that, if $K$ is nonreal, then every quadratic form over $K$ is a torsion form, and hence in this case the definition can be simplified to
$u(K)\,=\,\mathsf{sup}\{\mathsf{dim}(\varphi)\mid\varphi\mbox{ anisotropic formover }K\},$
which corresponds to how the $u$-invariant was originally defined by Lenz.
Let $v$ in the sequel be a valuation on $K$ with $\mathsf{char}(Kv)\neq 2$.
@@
We have $u(K)\geqslant[vK:2vK]\cdot u(Kv)$, except possibly when $K$ is real, $Kv$ is nonreal and $u(Kv)$ is odd. In any case, $u(K)\geqslant[vK:2vK]\cdot 2\lfloor\mbox{$\frac{u(Kv)}{2}$}\rfloor$.
@@
See [5, Theorem 5.2]. ∎
The following statement and its argument are essentially known; see e.g. [27, Theorem 4].
@@
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$. Then
@@
In view of 3.10 $(a)$, the hypothesis implies that $K$ is real if and only if $Kv$ is real. In particular, 4.1 yields that
@@
For this, we set $r=[vK:2vK]$ and $u=u(Kv)$ and assume without loss of generality that $r<\infty$. Note that $v^{-1}(2K)={\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. We fix representatives $c_{1},\dots,c_{r}\in{K}^{\times}$ of the different classes of ${K}^{\times}/{\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. Let $\varphi$ now be an anisotropic torsion form over $K$. We write $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$ where $\varphi_{1},\dots,\varphi_{r}$ are $v$-unimodular forms over $K$. Then $\overline{\varphi}_{1}^{v},\dots,\overline{\varphi}_{r}^{v}$ are anisotropic torsion forms over $K$, by 3.11. Hence $\mathsf{dim}(\varphi_{i})=\mathsf{dim}(\overline{\varphi}_{i}^{v})\leqslant u$ for $1\leqslant i\leqslant r$, whereby $\mathsf{dim}(\varphi)=\sum_{i=1}^{r}\mathsf{dim}(\varphi_{i})\leqslant r\cdot u$. This shows that $u(K)\leqslant r\cdot u$. ∎

### v2

@@ 3. $(c)$
$(b)$ If $\varphi$ is isotropic, then so is $\overline{\varphi}^{v}$, by 3.8. If $\overline{\varphi}$ is isotropic, then $\overline{\varphi}^{\prime}=\langle\overline{a}_{2},\dots,\overline{a}_{n}\rangle_{Kv}$ represents $\overline{a}_{1}$, and it follows by $(a)$ that $\varphi^{\prime}$ represents $a_{1}$, whereby $\varphi$ is isotropic.
$(c)$ We prove the equivalence by induction on $n$. If $n=0$, then $\varphi$ and $\overline{\varphi}^{v}$ are trivially hyperbolic. Assume that $n>0$. If $\varphi$ is anisotropic, then so is $\overline{\varphi}^{\prime}$, by $(b)$, and as $n>0$, it follows that neither $\varphi$ nor $\overline{\varphi}$ is hyperbolic. Assume that $\varphi$ is isotropic. By 3.5 $(b)$, there exist $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$ such that $\varphi\simeq\mathbb{H}_{K}\perp\langle b_{3},\dots,b_{n}\rangle_{K}$ and $\overline{\varphi}\simeq\mathbb{H}_{Kv}\perp\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$. Now, by Witt’s Cancellation Theorem [21, Chap. I, Theorem 4.2], $\varphi$ is hyperbolic if and only if $\rho=\langle b_{3},\dots,b_{n}\rangle_{K}$ is hyperbolic, and $\overline{\varphi}^{v}$ is hyperbolic if and only if $\overline{\rho}^{v}=\langle\overline{b}_{3},\dots,\overline{b}_{n}\rangle_{Kv}$ is hyperbolic. By the induction hypothesis, $\rho$ is hyperbolic if and only if $\overline{\rho}^{v}$ is hyperbolic. From this together we conclude that $\varphi$ is hyperbolic if and only if $\overline{\varphi}^{v}$ is hyperbolic. ∎
@@
Part of the following statement recovers [10, Theorem 1] and its corollary.
@@
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$. Let $\varphi$ be a quadratic form over $K$. Let $r\in\mathbb{N}$, $c_{1},\dots,c_{r}\in{K}^{\times}$ with $v(c_{i})\not\equiv v(c_{j})\,\,\mathsf{mod}\,\,2vK$ for $1\leqslant i<j\leqslant r$ and let $\varphi_{1},\dots,\varphi_{r}$ be $v$-unimodular quadratic forms over $K$ such that we have $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$. Then $\varphi$ is anisotropic, hyperbolic or torsion if and only if all residue forms $\overline{\varphi}_{1}^{v},\dots,\overline{\varphi}_{r}^{v}$ have the corresponding property.
@@
Preceding treatments of quadratic forms over complete nondyadically discretely valued (or more generally $2$-henselian nondyadically valued) fields often focussed more on describing the structure of the Witt ring of quadratic forms than on determining isotropy etc.  of individual quadratic forms. While we do not use the Witt ring in the present article, we wish to mention the following direct consequence of 3.11, which relates to the Witt ring point of view.
Two regular quadratic forms $\varphi$ and $\psi$ over a field of characteristic different from $2$ are called Witt equivalent if their anisotropic parts are isometric, which is if and only if $\varphi\perp-\psi$ is hyperbolic.
Corollary 3.12.
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$. Let $r\in\mathbb{N}$, $c_{1},\dots,c_{r}\in{K}^{\times}$ with $v(c_{i})\not\equiv v(c_{j})\,\,\mathsf{mod}\,\,2vK$ for $1\leqslant i<j\leqslant n$ and let $\vartheta_{1},\psi_{1},\dots,\vartheta_{r},\psi_{r}$ be $v$-unimodular quadratic forms over $K$. Then the forms $c_{1}\vartheta_{1}\perp\ldots\perp c_{r}\vartheta_{r}$ and $c_{1}\psi_{1}\perp\ldots\perp c_{r}\psi_{r}$ over $K$ are Witt equivalent if and only if, for $1\leqslant i\leqslant r$, the forms $\overline{\vartheta}_{i}$ and $\overline{\psi}_{i}$ over $Kv$ are Witt equivalent.
Proof.
Letting $\varphi_{i}=\vartheta_{i}\perp-\psi_{i}$ for $1\leqslant i\leqslant r$, the statement follows from the criterion for the hyperbolicity of the quadratic form $\varphi=c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$ over $K$ which is contained in 3.11. ∎
@@
When restricting attention to complete discrete valuation rings and their fraction fields, most (if not all) of the results in this section are well-known. To the best of our knowledge these first appeared explicitly in Durfee’s doctoral research [10] for fields with a complete rank-$1$ valuation. Inspection reveals that Durfee’s proofs (and other versions found later) mostly work without changes for general valued fields $(K,v)$ with $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ and $2\in\mathcal{O}_{v}^{\times}$, thus in particular for non-dyadic henselian valued fields, which would be sufficient for the purposes in the rest of this article. For our own and the reader’s convenience, we have included here an elementary and self-contained exposition which postpones adding the assumption $1+\mathfrak{m}_{v}\subseteq K^{\times 2}$ until the very end.
@@ 4. The generalised $u$-invariant
By the Artin-Schreier Theorem, a field $K$ admits a field ordering if and only if 1-1 is not a sum of squares in $K$; see e.g. [21, Chap. VIII, Theorem 1.10]. This motivates to call the field $K$ real in this case, and nonreal otherwise. Given a field ordering, a quadratic form is assigned a corresponding signature, and there is a notion of being positive or negative definite at this ordering. A quadratic form which is positive definite at some ordering cannot be universal, since it cannot represent 1-1. Therefore the definition of the $u$-invariant formulated in the introduction for nonreal fields would have value $\infty$ for any real field. Elman and Lam [12] suggested a variation to the definition of the $u$-invariant that is meaningful also for real fields while not changing its meaning for nonreal fields. We refer to [32, Chap. 8] for a general discussion of the $u$-invariant, including several variations on how to define it for real fields on the one hand and for fields of characteristic $2$ on the other hand.
In this section, we discuss relations between the values of the $u$-invariant for a given valued field and its residue field. There seems a want in the literature for such a treatment covering both real fields and fields of characteristic $2$ in a compatible way. Most of the statements here are well-known in the case where the valued field is nonreal and the residue field is of characteristic different from $2$. Including the case of real fields in the discussion leads to some technicalities, which can be skipped if one restricts the attention to nonreal fields.
Let $K$ be a field. Recall that, if $\mathsf{char}(K)\neq 2$, we defined a quadratic form $\varphi$ over $K$ to be a torsion form if $m\times\varphi=\varphi\perp\ldots\perp\varphi$ is hyperbolic for some $m\in\mathbb{N}^{+}$. When $\mathsf{char}(K)=2$, we take the convention that every quadratic form over $K$ is torsion. The general $u$-invariant of $K$ as defined by Elman and Lam in [12] is given by
@@
Note that, if $K$ is nonreal, then every anisotropic quadratic form over $K$ is a torsion form, and hence in this case the definition can be simplified to
$u(K)\,=\,\mathsf{sup}\{\mathsf{dim}(\varphi)\mid\varphi\mbox{ anisotropic quadratic form over }K\},$
which corresponds to how the $u$-invariant was originally defined by Lenz. This further agrees with the definition mentioned in the introduction when $K$ is nonreal of characteristic different from $2$.
Pfister’s Local-Global Principle [21, Chap. VIII, Sect. 3] states that, over a real field, a regular quadratic form is torsion if and only if it has signature zero at every field ordering. Therefore, if $K$ is real, then the $u$-invariant of $K$ has to be even or infinite.
The following construction will be used to produce certain torsion forms in characteristic different from $2$.
Lemma 4.1.
Assume that $\mathsf{char}(K)\neq 2$. Let $\vartheta$ be a regular quadratic form over $K$. Let $c\in{K}^{\times}$ and $k\in\mathbb{N}$ be such that $c$ is a sum of $2^{k}$ squares in $K$. Then $\vartheta\perp-c\vartheta$ is a torsion form. More precisely, $2^{k}\times(\vartheta\perp-c\vartheta)$ is hyperbolic.
Proof.
The hypothesis implies that the form $2^{k}\times\langle 1,-c\rangle_{K}$ is isotropic, and since it is a Pfister form, it follows by [21, Theorem X.2.9] that it is hyperbolic. Consequently, $2^{k}\times\langle a,-ac\rangle$ is hyperbolic for every $a\in{K}^{\times}$. Let $n=\mathsf{dim}(\vartheta)$. By 3.2, there exist $a_{1},\dots,a_{n}\in{K}^{\times}$ such that $\vartheta\simeq\langle a_{1},\dots,a_{n}\rangle_{K}$. Then 2k×(ϑcϑ)
i=1n
(2k×ai,caiK)
[…]
, which is hyperbolic. ∎
Let $v$ in the sequel be a valuation on $K$.
Lemma 4.2.
Let $n\in\mathbb{N}$ and let $\psi$ be an $n$-dimensional quadratic form over $Kv$. Then there exists an $n$-dimensional quadratic form $\Psi$ over $\mathcal{O}_{v}$ which reduces modulo $\mathfrak{m}_{v}$ to $\psi$. Moreover, if $\psi$ is anisotropic, then any such form $\Psi$ is anisotropic over $K$ and satisfies $v(\Psi(x))\in 2vK$ for every $x\in K^{n}\smallsetminus\{0\}$.
Proof.
See [2, Proposition 4.2]. ∎
In some cases where we will use 4.2, we want the lifted form $\Psi$ in addition to be a torsion form over $K$. To achieve this some extra considerations are necessary when $K$ is real.
For a ring $R$ we denote by $\mathsf{\Sigma}{R}^{(2)}$} the set of elements of $R$ which are sums of squares in $R$.
Lemma 4.3.
Assume that $Kv$ is nonreal and $\mathsf{char}(K)\neq 2$. The following hold:
  1. $(a)$
    Every element of $Kv$ is the residue modulo $\mathfrak{m}_{v}$ of an element of $\mathcal{O}_{v}\cap{\mathsf{\Sigma}{K}^{(2)}}$.
  2. $(b)$
    Every anisotropic even-dimensional quadratic form over $Kv$ is the reduction modulo $\mathfrak{m}_{v}$ of a quadratic form defined over $\mathcal{O}_{v}$ which is an anisotropic torsion form over $K$.
Proof.
$(a)$ If $\mathsf{char}(Kv)=2$, then for $x\in\mathcal{O}_{v}$ we have $x+2(x^{2}+1)\equiv x\,\,\mathsf{mod}\,\,\mathfrak{m}_{v}$ and $x+2(x^{2}+1)=2(x+\frac{1}{4})^{2}+\frac{15}{8}\in{\mathsf{\Sigma}{K}^{(2)}}$. If $Kv$ is nonreal with $\mathsf{char}(Kv)\neq 2$, then there exists $z\in{\mathsf{\Sigma}{\mathcal{O}}^{(2)}}$ with $z\equiv-1\,\,\mathsf{mod}\,\,\mathfrak{m}_{v}$, and for $x\in\mathcal{O}_{v}$ we obtain that $\frac{1}{4}((x+1)^{2}+z(x-1)^{2})\in{\mathsf{\Sigma}{\mathcal{O}_{v}}^{(2)}}$ and $x\equiv\frac{1}{4}((x+1)^{2}+z(x-1)^{2})\,\,\mathsf{mod}\,\,\mathfrak{m}_{v}$.
$(b)$ Consider an arbitrary anisotropic $2$-dimensional form $\beta$ over $Kv$. By $(a)$, we may choose $a,b,c\in\mathcal{O}_{v}\cap{\mathsf{\Sigma}{K}^{(2)}}$ such that the $2$-dimensional quadratic form $B=aX_{1}^{2}+bX_{1}X_{2}-cX_{2}^{2}$ over $\mathcal{O}_{v}$ reduces modulo $\mathfrak{m}_{v}$ to $\beta$. Since $\beta$ is anisotropic over $Kv$, so is $B$ over $K$. Note that $B\simeq a\langle 1,-(b^{2}+4ac)\rangle$ over $K$ and $b^{2}+4ac\in{\mathsf{\Sigma}{K}^{(2)}}$, whence it follows by 4.1 that $B$ is a torsion form. This shows the claim for $2$-dimensional quadratic forms. By [11, Propositions 7.29 and 7.31] every even-dimensional quadratic form is an orthogonal sum of $2$-dimensional quadratic forms. Hence the statement follows in general. ∎
The following is a partial extension of [5, Theorem 5.2], including now the case of residue characteristic $2$.
@@ 2. $(b)$
We have
$u(K)\geqslant[vK:2vK]\cdot 2\lfloor\mbox{$\frac{u(Kv)}{2}$}\rfloor\,.$
Moreover,
$u(K)\geqslant[vK:2vK]\cdot u(Kv)$
holds under any of the following conditions: (i)(i)$K$ is nonreal; (ii)(ii)KvKv is real; $(iii)$  $u(Kv)$ is even; $(iv)$ $v({\mathsf{\Sigma}{K}^{(2)}}\smallsetminus\{0\})\not\subseteq 2vK$.
@@
In the case where $Kv$ is real, the statement is fully covered by [5, Theorem 5.2]. Hence we only consider the complementary situation and assume from now on that $Kv$ is nonreal.
Let $n,r\in\mathbb{N}$ arbitrary with $2^{r}\leqslant[vK:2vK]$ and $n\leqslant u(Kv)$, and assume that $n$ is even in the case where $K$ is real and $v({\mathsf{\Sigma}{K}^{(2)}}\smallsetminus\{0\})\subseteq 2vK$. We will show that $u(K)\geqslant 2^{r}n$. This will establish the statement.
By the choice of $n$ there exists an anisotropic $n$-dimensional quadratic form $\varphi$ over $Kv$. Then $\varphi$ is the reduction modulo $\mathfrak{m}_{v}$ of an $n$-dimensional quadratic form $\Phi$ over $\mathcal{O}_{v}$, and by 4.3 we may choose $\Phi$ to be a torsion form over $K$ when $n$ is even. Also if $K$ is nonreal, $\Phi$ is torsion. In any case, since $\varphi$ is anisotropic, it follows by 4.2 that $v(\Phi(x))\in 2vK$ for every $x\in K^{n}\smallsetminus\{0\}$.
We may assume that $r\geqslant 1$ unless $vK=2vK$. By the choice of $r$, we may choose $c_{1},\ldots,c_{r}\in{K}^{\times}$ such that $v(c_{1}),\dots,v(c_{r})$ represent $\mathbb{F}_{2}$-linearly independent classes in $vK/2vK$, and we may choose to have $c_{1}\in-{\mathsf{\Sigma}{K}^{(2)}}$ except when $v({\mathsf{\Sigma}{K}^{(2)}}\smallsetminus\{0\})\subseteq 2vK$. We denote by $\mathcal{P}$ the power set of $\{1,\dots,r\}$, and for $I\in\mathcal{P}$ we set $c_{I}=\prod_{i\in I}c_{i}$. We consider the quadratic form Ψ=
[…]
cIΦ
[…]
over $K$. It follows by the properties of $\Phi$ and $c_{1},\dots,c_{r}$ that $\Psi$ is anisotropic. We claim that $\Psi$ is torsion. This is obvious when $\Phi$ is torsion. Assume that $\Phi$ is not torsion. In this case $n$ is odd, $vK\neq 2vK$, $r\geqslant 1$ and $c_{1}\in{\mathsf{\Sigma}{K}^{(2)}}$, and since $\Psi=\Theta\perp c_{1}\Theta$ with Θ=
[…]
cIΦ
[…]
where $\mathcal{P}^{\prime}$ is the power set of $\{2,\dots,r\}$, we conclude by 4.1 that $\Psi$ is torsion. Hence $u(K)\geqslant\mathsf{dim}(\Psi)=2^{r}n$. Having this for all $n$ and $r$ as above, the statement follows. ∎
We obtain the following lower bound for the growth of the $u$-invariant under finitely generated field extensions in terms of the transcendence degree. The argument is well-known at least for diagonal quadratic forms over nonreal fields, see e.g. [35, Cor. 2 $(ii)$].
Corollary 4.5.
Let $F/K$ be a finitely generated field extension and let $n$ be its transcendence degree. Then there exists a finite field extension $L/K$ such that $u(F)\geqslant 2^{n}u(L)$.
Proof.
The statement is trivial when $n=0$. Consider the case where $n=1$. In this case $F/K$ is a function field in one variable. We pick any discrete valuation $v$ on $F$ which is trivial on $K$. We set $L=Fv$. Then $L/F$ is a finite field extension and $\mathbb{Z}:2\mathbb{Z}]=2$. Moreover, if $F$ is real but $Fv$ is nonreal, then by [4, Proposition 4.2], there exists $x\in{\mathsf{\Sigma}{F}^{(2)}}$ with $v(x)=1$. Hence, in any case we obtain by 4.4 that $u(F)\geqslant 2u(L)$. This proves the statement for $n=1$. The general statement follows by induction on $n$. ∎
Determining the $u$-invariant of a valued field in terms of the value group and the $u$-invariant of the residue field is possible in residue characteristic different from $2$ provided that roots of quadratic equations can be lifted from the residue field. This is also well-known over nonreal fields, see e.g. [35, Theorem 4].
@@
Assume that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$ and $\mathsf{char}(Kv)\neq 2$. Then
@@
In view of 3.10 $(a)$, the hypothesis implies that $K$ is real if and only if $Kv$ is real. In particular, 4.4 yields that
@@
For this, we set $s=[vK:2vK]$ and $u=u(Kv)$ and assume without loss of generality that $s<\infty$. Note that $v^{-1}(2K)={\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. We fix representatives $c_{1},\dots,c_{s}\in{K}^{\times}$ of the different classes of ${K}^{\times}/{\mathcal{O}}^{\times}_{v}{K}^{\times 2}$. Let $\varphi$ now be an anisotropic torsion form over $K$. We write $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{s}\varphi_{s}$ where $\varphi_{1},\dots,\varphi_{s}$ are $v$-unimodular forms over $K$. Then $\overline{\varphi}_{1}^{v},\dots,\overline{\varphi}_{s}^{v}$ are anisotropic torsion forms over $K$, by 3.11. Hence $\mathsf{dim}(\varphi_{i})=\mathsf{dim}(\overline{\varphi}_{i}^{v})\leqslant u$ for $1\leqslant i\leqslant s$, whereby $\mathsf{dim}(\varphi)=\sum_{i=1}^{s}\mathsf{dim}(\varphi_{i})\leqslant s\cdot u$. This shows that $u(K)\leqslant s\cdot u$. ∎
A similarly general statement covering the case of residue characteristic $2$ seems out of reach. Only the following statement is known to hold without restriction on the residue characteristic. One may trace it back to Kaplansky [18], who observes that $u(K)=2u(k)$ in the case where K=k((t))K=k(\!(t)\!) for a field $k$ with $\mathsf{char}(k)\neq 2$.
If $\mathsf{char}(K)=2$, we denote by $K:K^{(2)}]$ the degree of imperfection of $K$, i.e. the degree of the field extension $K/K^{(2)}$ where $K^{(2)}=\{x^{2}\mid x\in K\}$.
Theorem 4.7 (Kaplansky; Mamone-Moresi-Wadsworth).
Assume that $v$ is discrete and henselian. If $\mathsf{char}(K)=2$, then assume further that $K:K^{(2)}]=2[Kv:Kv^{(2)}]$. Then $u(K)=2u(Kv)$.
Proof.
When $\mathsf{char}(Kv)\neq 2$ this follows from 4.6. See [25, Theorem 2] for the case when $\mathsf{char}(Kv)=2$. ∎
Note that, if $\mathsf{char}(K)=2$ and $v$ is complete and discrete, then $K:K^{(2)}]=2[Kv:Kv^{(2)}]$ holds, whereby 4.7 applies. In a more general context, when $\mathsf{char}(Kv)=2$, one may have an inequality $u(K)>[vK:2vK]\cdot u(Kv)$, even when $(K,v)$ is henselian and complete. We sketch an example, where we omit the verification of valuation-theoretic details.
Example 4.8.
Let $\overline{\mathbb{F}}_{2}$ denote the algebraic closure of $\mathbb{F}_{2}$ and $K$ the perfect closure (maximal purely inseparable extension) of $\overline{\mathbb{F}}_{2}(t)$ or of $\overline{\mathbb{F}}_{2}(\!(t)\!)$. That is, $K=\overline{\mathbb{F}}_{2}(t^{2^{-n}}|\,n\in\mathbb{N})$ or $K=\overline{\mathbb{F}}_{2}(\!(t)\!)(t^{2^{-n}}|\,n\in\mathbb{N})$, respectively. Let $v$ denote the unique extension of the $t$-adic valuation on $\overline{\mathbb{F}}_{2}(t)$ and $\overline{\mathbb{F}}_{2}(\!(t)\!)$ to $K$. We also denote the natural extension of this valuation to the completion $K^{v}$ by $v$. Note that $(K^{v},v)$ is a complete rank-$1$ valued field. In particular, it is henselian. Since $K$ is perfect, so is $K^{v}$, whereby trivially $1+\mathfrak{m}_{v}\subseteq{(K^{v})}^{\times 2}$. The value group $vK$ is $2$-divisible, and the residue field $Kv=\overline{\mathbb{F}}_{2}$ is algebraically closed, whereby $[vK:2vK]=1=u(Kv)$. But $u(K)>1$, because the quadratic form $X_{1}^{2}+X_{1}X_{2}+t^{-1}X_{2}^{2}$ is anisotropic over $K^{v}$.
With a bit more work, an example can also be constructed of a complete henselian valued field $(K,v)$ with $\mathsf{char}(K)=0$, $\mathsf{char}(Kv)=2$, and $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$, whereas $u(K)>[vK:2vK]\cdot u(Kv)$.